# Graph-level Representation Learning with Joint-Embedding Predictive Architectures

### v1

@@ 3.1 Spatial Partitioning
We base the initial part of the Graph-JEPA architecture on the recent work of He et al. (2023), but similar ideas consisting of graph partitioning have been proposed before for Graph SSL (Jin et al., 2020). This step consists of creating different subgraphs (patches) of a graph, similar to how Vision Transformers (ViT) (Dosovitskiy et al., 2020) operate on images. We rely on the METIS (Karypis & Kumar, 1998) graph clustering algorithm, which partitions a graph into a pre-defined, non-overlapping number of clusters $p$ (Figure 2a), such that the number of within-cluster links is much higher than between-cluster links. Note that having non-overlapping subgraphs can be problematic since edges can be lost in this procedure and it is possible to end up with empty ”subgraphs”. We would also like to preserve the notion of locality in each subgraph while relating it to others that are close in terms of shared nodes. To create this overlap and avoid completely empty subgraphs, a one-hop neighborhood expansion of all the nodes belonging to a subgraph is performed (Figure 2b).
@@ 3.2 Subgraph Embedding
After partitioning the graph, we learn a representation for each subgraph through a GNN (Figure 2c.). The specific choice of GNN architecture is arbitrary and depends on what properties we wish to induce in the representation. The learned node embeddings are mean pooled to create a vector representation for each subgraph: $\{h_{1}...h_{p}\},h\in\mathbb{R}^{d}$. Given that these embeddings will be used as context or target variables, it is important to provide additional information regarding the subgraphs to help guide the predictor network. Thus, we propose to use a positional embedding for each subgraph, which is implemented as the maximum Random Walk Structural Embedding (RWSE) of all the nodes in that subgraph. In this way, the position is characterized in a global and consistent manner for each patch. Formally, a RWSE (Dwivedi et al., 2021) for a node $v$ can be defined as:
$P_{v}=(M_{ii},M^{2}_{ii},\ldots,M^{k}_{ii}),P_{v}\in\mathbb{R}^{k},~{}~{}~{}~{}~{}M^{k}=(D^{-1}A)^{k}$ (1)
where $M^{k}$ is the random-walk diffusion matrix of order $k$ and $i$ is the index of node $v$ in the adjacency matrix. Therefore, $M^{k}_{ii}$ encodes the probability of node $v$ landing to itself after a $k$-step random walk. Given a subgraph $l$, let $V_{l}$ denote the set of nodes in $l$. The subgraph RWSE is then defined as:
@@ 3.3 Context and Target Encoding
Given the subgraph representations and their respective positional embeddings, we frame the Graph-JEPA prediction task in a similar manner to I-JEPA (Assran et al., 2023). The goal of the network is to predict the latent embeddings of randomly chosen target subgraphs, given one random context subgraph. The prediction is conditioned on positional information regarding each subgraph. At each training step, we choose one random subgraph as the context $x$ and $m$ others as targets $Y=\{y_{1},\ldots,y_{m}\}$ (Figure 2d). These subgraphs are processed by the context and target encoders (Figure 2e) which are parametrized by Transformer encoder blocks (without self-attention for the context) where normalization is applied at first (Xiong et al., 2020). The target encoder uses Hadamard self-attention (He et al., 2023), but other choices, such as the standard self-attention mechanism (Vaswani et al., 2017) are perfectly viable. We can summarize this step as:
$z^{x}=E_{context}(x),z^{x}\in\mathbb{R}^{d},~{}~{}~{}Z^{y}=E_{target}(Y),Z^{y}\in\mathbb{R}^{m\times d}$ (3)
At this stage, we could use the predictor network to directly predict $Z^{y}$ from $z^{x}$. This is the typical formulation of JEPAs, also followed by Assran et al. (2023). We argue that learning how to organize concepts for abstract objects such as graphs or networks directly in Euclidean space is suboptimal. In the following section, we propose a simple trick to bypass this problem using the encoding and prediction mechanisms in Graph-JEPA. The following subsections and the ablations studies in Section 4.3 will provide additional insights.
@@ 3.4 Latent Target Prediction
Different works in Neuroscience have shown the importance of learning hierarchically consistent concepts (Deco et al., 2021), especially during infancy and young age (Rosenberg & Feigenson, 2013). Networks in the real world often conform to some concept of hierarchy (Moutsinas et al., 2021) and this assumption is frequently used when learning graph-level representations (Ying et al., 2018). Thus, we conjecture that Graph-JEPA should operate in a hyperbolic space, where learned embeddings implicitly organize hierarchical concepts(Nickel & Kiela, 2017; Zhao et al., 2023). This gives rise to another issue: Hyperbolic (Poincarè) embeddings are known to have several tradeoffs related to dimensionality (Sala et al., 2018; Guo et al., 2022), which severely limits the expressive ability of the model. Given that graphs can describe very abstract concepts, high expressivity in terms of model parameters is preferred. In summary, we would ideally like to have a high-dimensional latent code that has a concept of hyperbolicity built into it.
To achieve this, we think of the target embedding as a high-dimensional representation of the hyperbolic angle, which allows us to describe each target patch through its position in the 2D unit hyperbola. Formally, given a target patch $l$, its embedding $Z^{y}_{l}$ and positional encoding $P_{l}$, we first express the latent target as:
$\alpha^{y}_{l}=\frac{1}{N}\sum_{n=1}^{d}{Z^{y}_{l}}_{n},~{}~{}~{}\psi^{y}_{l}=\begin{pmatrix}\;cosh(\alpha^{y}_{l})\;\\ \\ \;sinh(\alpha^{y}_{l})\;\end{pmatrix}$ (4)
where $cosh(\cdot)$ and $sinh(\cdot)$ are the hyperbolic cosine and sine functions respectively. The predictor module is then tasked with directly locating the target in the unit hyperbola given the context embedding and the target patch’s positional encoding.
$\hat{\psi}^{y}_{l}=MLP(LayerNorm(z^{x}+P_{l})),\hat{\psi}^{y}_{l}\in\mathbb{R}^{2}$ (5)
This allows us to frame the learning procedure as a simple regression problem and the whole network can be trained end-to-end (Figure 2f). In practice, we use the smooth L1 loss as the distance function, as it is less sensitive to outliers compared to the L2 loss (Girshick, 2015):
@@
Thus, we are effectively measuring how far away the context and target patches are in the unit hyperbola of the plane, but the targets are actually described through a high dimensional latent code (Eq. 4). We explicitly show the differences between this choice and using the Euclidean or Hyperbolic distances as energy functions (in the latent space) in Section 4.3. Our proposed pretraining objective forces the context encoder to understand the differences in the hyperbolic angle between the target patches, which can be thought of as establishing a hierarchy.
An asymmetric design of the predictor and target networks (in terms of parameters) is used since it is reported to prevent representation collapse in self-supervised contrastive techniques (Chen et al., 2020; Baevski et al., 2022). We also utilize stop-gradient for the target encoder and update its weights using an Exponential Moving Average of the context encoder, as done in other works that utilize JEPAs (Assran et al., 2023; Grill et al., 2020; Chen & He, 2021). In the next section, we provide insights into these choices and why they are important. Finally, to produce the unique graph-level representation when fine-tuning, we simply feed all the subgraphs through the trained target encoder and then use mean pooling, obtaining a single feature vector $z_{G}\in\mathbb{R}^{d}$ that represents the whole graph.
3.5 Why does Graph-JEPA work?
In this subsection, we aim to provide a discussion regarding the design choices for Graph-JEPA and also why our formulation makes sense in the graph domain. We make the following assumptions for our analysis: i) The predictor network is linear; ii) We consider the encoded context features $z^{x}$ and target coordinates $\psi^{y}_{l}$. Without loss of generality, we assume that there is a one-to-one correspondence between context and target patches. (This holds in practice due to Eq. 5; iii) The problem is treated as a least-squares problem in finite-dimensional vector spaces over the field of reals $\mathbb{R}$. Based on our assumptions, we can rewrite the context features as $X\in\mathbb{R}^{n\times d}$, the target coordinates as $Y\in\mathbb{R}^{n\times 2}$, and the weights of the linear model as $W\in\mathbb{R}^{d\times 2}$. The objective of the predictor is:
@@ 3.5 Why does Graph-JEPA work?
By plugging Eq. 8 into Eq. 7 and rewriting the difference by factorizing $Y$ we get:
$\operatorname*{arg\,min}\limits_{W}\left\|(X(X^{T}X)^{-1}X^{T}-I_{n})Y\right\|^{2}$ (9)
Thus, to find an optimal (linear) predictor we must find the orthogonal projection of $Y$ onto the orthogonal complement of a subspace of $Col(X)$. This simply translates into finding the linear combination that is closest (in terms of $\left\|\cdot\right\|^{2}$) to $Y$ in a particular subspace. Similarly to what was shown in Richemond et al. (2023), we argue that this provides a key intuition behind the JEPA design. The target encoder which produces $Y$ must not share weights or be optimized with the context encoder. If that were the case, the easiest solution to Eq. 9 would be using a vector that is orthogonal to itself, i.e., the $\mathbf{0}$ vector, leading to representation collapse. Note that in practice, even though the predictor network doesn’t have to be linear, it would still fail to perfectly learn a linear relationship (as shown above). This would render the architecture suboptimal.
We now argue about the existence of hyperbolicity due to our latent prediction task and the necessity of the asymmetric predictor. The predictor network is linearly approximating the behavior of the unit hyperbola such that it best matches with the generated target coordinates (Eq 6). Thus, the network is actually trying to estimate a space that can be considered a particular section of the hyperboloid model (Reynolds, 1993). In the hyperboloid model, hyperbolas appear as hyperbolic geodesics. We are therefore evaluating our energy in a (very) restricted part of hyperbolic space. As mentioned before, we find this task to offer great flexibility as it is straightforward to implement, works well in practice (Section 4.2), and is computationally efficient compared to the hyperbolic distance used to typically learn hyperbolic embeddings (Nickel & Kiela, 2017). In our current setting, this approximation cannot perfectly capture the dynamics of the unit hyperbola, given the linearity of the predictor. In practice, with a non-linear predictor network, we might immediately run into a degenerate solution where the target encoder and predictor output collapse representations by immediately minimizing the training objective. It is therefore important to parametrize the predictor using a simpler network that is less expressive but can capture the correct dynamics over the training process.
@@ 4 Experiments
The experimental section introduces the empirical evaluation of the Graph-JEPA model in terms of downstream performance on different graph datasets and tasks, as well as various ablation studies that provide more insights into our design choices.
@@ 4.1 Experimental setting
We use the TUD datasets (Morris et al., 2020) as commonly done for graph-level SSL (Suresh et al., 2021; Tan et al., 2023). We utilize seven different graph-classification datasets: PROTEINS, MUTAG, DD, REDDIT-BINARY, REDDIT-MULTI-5K, IMDB-BINARY and IMDB-MULTI. For all classification experiments, we report the accuracy over five runs (with different seeds) of ten-fold cross validation. It is worth noting that we retrain the Graph-JEPA model for each fold, without ever having access to the testing partition both in the pretraining and fine-tuning stage. For graph regression, we use the ZINC dataset and report the Mean Squared Error over ten runs (with different seeds), given that the testing partition is already provided. More details regarding the training procedure are available in Appendix A.1.
4.2 Downstream performance
For the experiments on downstream performance, we follow Suresh et al. (2021) and also report the results of a fully supervised Graph Isomorphism Network (GIN) (Xu et al., 2018), denoted F-GIN. As can be seen in Table 1, Graph-JEPA achieves competitive results on all datasets, setting the state-of-the-art as a pretrained backbone on five different datasets and coming second on one. Overall, our proposed framework learns semantic embeddings that work well on different types of graphs, showing that Graph-JEPA can indeed be utilized as a general pretraining method for graph-level SSL. Notably, Graph-JEPA works well for both classification and regression and performs better than F-GIN on all classification datasets.
We further explore the performance of our model on the synthetic EXP dataset (Abboud et al., 2020). The goal behind this experiment is to empirically verify if Graph-JEPA can learn highly expressive graph representations, in terms of graph isomorphisms. The results in Table 2 show that our model is able to perform much better than commonly used GNNs. Given its local and global exchange of information, this result is to be expected. Most importantly, Graph-JEPA closely rivals the flawless performance of Graph-MLP-Mixer (He et al., 2023), which is trained in a supervised manner.
Table 1: Performance of different graph SSL techniques ordered by pretraining type. The results of the competitors are taken as the best values from (Hassani & Khasahmadi, 2020; Suresh et al., 2021; Tan et al., 2023). ”-” indicates missing values from the literature. The best results are reported in boldface, the second best in italics.
F-GIN (Xu et al., 2018) 72.39 $\pm$ 2.76 90.41 $\pm$ 4.61 74.87 $\pm$ 3.56 86.79 $\pm$ 2.04 53.28 $\pm$ 3.17 71.83 $\pm$ 1.93 48.46 $\pm$ 2.31 0.254 $\pm$ 0.005
RU-GIN (Xu et al., 2018) 69.03 $\pm$ 0.33 87.61 $\pm$ 0.39 74.22 $\pm$ 0.30 58.97 $\pm$ 0.13 27.52 $\pm$ 0.61 51.86 $\pm$ 0.33 32.81 $\pm$ 0.57 0.809 $\pm$ 0.022
AD-GCL-OPT (Suresh et al., 2021) 73.81 $\pm$ 0.46 89.70 $\pm$ 1.03 75.10 $\pm$ 0.39 85.52 $\pm$ 0.79 54.93 $\pm$ 0.43 72.33 $\pm$ 0.56 49.89 $\pm$ 0.66 0.544 $\pm$ 0.004
GraphMAE (Hou et al., 2022) 75.30 $\pm$ 0.39 88.19 $\pm$ 1.26 - 88.01 $\pm$ 0.19 - 75.52 $\pm$ 0.66 51.63 $\pm$ 0.52 -
Graph-JEPA 75.67 $\pm$ 3.78 91.25 $\pm$ 5.75 78.64 $\pm$ 2.35 91.99 $\pm$ 1.59 56.73 $\pm$ 1.96 73.68 $\pm$ 3.24 50.69 $\pm$ 2.91 0.434 $\pm$ 0.014
Table 2: Classification accuracy on the synthetic EXP dataset (Abboud et al., 2020), which contains 600 pairs of non-isomorphic graphs that are indistinguishable by the 1-WL test. The competitor models are proposed in (Kipf & Welling, 2016a; Bresson & Laurent, 2017; Xu et al., 2018; Dwivedi & Bresson, 2020; He et al., 2023)
. Model GCN GatedGCN GINE GraphTransformer Graph-MLP-Mixer Graph-JEPA Accuracy 51.90 $\pm$ 1.96 51.73 $\pm$ 1.65 50.69 $\pm$ 1.39 52.35 $\pm$ 2.32 100.00 $\pm$ 0.00 98.77 $\pm$ 0.99
4.3 Exploring the Graph-JEPA latent space
As discussed in Section 3.4, the choice of energy function has a big impact on the learned representations. We argued against the use of simple Euclidean or Poincarè embeddings previously and now provide empirical evidence for this conjecture in Table 3. Our results reveal that learning the distance between patches in the 2D unit hyperbola provides a simple way to get the advantages of both embedding types. Hyperbolic embeddings must be learned in lower dimensions due to stability issues (Yu & De Sa, 2021), while Euclidean ones do not properly reflect the dependencies between subgraphs. Our results suggest that the hyperbolic distance seems to be a better choice than the Euclidean distance in lower dimensions, but it is very unstable and expensive computationally in high dimensions. The proposed approach provides the best overall results, with the Euclidean distance performing negligibly better in a single case. We provide a qualitative example that describes how the embedding space is altered when using our proposed latent objective in Appendix A.3.
Table 3: Comparison of Graph-JEPA performance for different distance functions. We used a latent space of dimension 128 for Poincarè embeddings due to optimization issues, which was still problematic for the IMDB-B dataset. LD stands for Lower Dimension, where we use a much smaller embedding size (32).
IMDB-B 73.68 $\pm$ 3.24 73.76 $\pm$ 3.46 NaN 72.5 $\pm$ 3.97 73.4 $\pm$ 4.07
4.4 Ablation studies
In the following, we present various ablation studies to get a better understanding of the different design choices of Graph-JEPA. For all ablations, we consider 4 out of the 8 datasets initially presented in Table 1, making sure to choose different graph types for a fair comparison. All results are summarized in Figure 3, but we also provide the results in tabular format in Appendix A.2 for additional clarity.
Positional embedding. Following He et al. (2023), it is possible to use the RWSE of the patches as conditioning information. Formally, let $B\in\{0,1\}^{p\times N}$ be the patch assignment matrix, such that $B_{ij}=1$ if $v_{j}\in p_{i}$. We can calculate a coarse patch adjacency matrix $A^{\prime}=BB^{T}\in\mathbb{R}^{p\times p}$, where each $A_{ij}^{\prime}$ contains the node overlap between $p_{i}$ and $p_{j}$. The RWSE can be calculated as described in Eq. 1 (considering $A^{\prime}$). We test Graph-JEPA with these relative positional embeddings and find that they still provide good performance, but consistently fall behind the node-level (global) RWSE that we employ in our formulation (Figure 3a). Indeed, a great issue for the relative embeddings is that the number of shared neighbors obscures the local peculiarities of each patch and this adds more variance to the results.
Using different Self-Attention mechanisms. As stated in Section 3.3, Graph-JEPA uses Hadamard self-attention (He et al., 2023), which provides a strong inductive bias for graphs. It is possible to make no assumptions about the data and render the prediction task more challenging, by using the standard self-attention mechanism (Vaswani et al., 2017). We show the results of this change in Figure 3b. The results reveal that the performance is slightly better but more unstable, i.e., additional variance in the results. This is to be expected since, as mentioned previously, using the standard self-attention mechanism emphasizes a lack of inductive bias regarding the input data.
Random subgraphs. A natural question that arises in our framework is how to design the spatial partitioning procedure. Using a structured approach like METIS is intuitive and leads to favorable results. Another option would be to simply extract random, non-empty subgraphs as context and target. As can be seen in Figure 3c, the random patches provide strong performance as well but do not show any improvement. Even though our results show that it might not be necessary to use a structured way to extract the patches, not all graphs are equally easy to sample randomly from. Thus, we advocate extracting subgraphs with METIS as it is a safer option in terms of generalizability across different graphs and what inductive biases it provides.
Figure 3: Ablation studies: a. Performance when using absolute (node-level) vs relative (patch-level) RWSEs. b. Results when using the standard self-attention mechanism vs a graph-specific one. c. Performance when using METIS subgraphs vs using random subgraphs. We remind the reader that the MAE is reported on ZINC and Accuracy on the other datasets. The error bars represent the average standard deviation values over the runs in order to report results similarly to the tables in previous sections. Best viewed in color.
5 Conclusion
In this work, we introduce the first Joint Embedding Predictive Architecture (JEPA) for graph-level Self-Supervised Learning (SSL). A proper design of the model both in terms of data preparation and pretraining objective reveals that it is possible for a neural network model to self-organize the semantic knowledge embedded in a graph, demonstrating competitive performance in both graph classification and regression. Future research directions include extending the method to node and edge-level learning, exploring the expressiveness of Graph-JEPA, and gaining insights into the optimal geometry of the embedding space for graph SSL.
Appendix A Appendix
A.1 Experimental setting
For the datasets that do not natively have node and edge features, we use a simple constant (0) initialization. The subgraph embedding GNN (Figure 2c) consists of the GIN operator with support for edge features (Hu et al., 2019) a.k.a GINE. For fine-tuning, we feed the high-dimensional representation of the hyperbolic angle produced by the target encoder to a linear model with L2 regularization. For the classification datasets, that corresponds to a Logistic Regression with L2 regularization, while for ZINC it corresponds to a Ridge Regression. We also provide details regarding the hyperparameters of the JEPA design for Graph-JEPA in Table 4. The neural network is trained using the Adam optimizer (Kingma & Ba, 2014). All experiments were performed on a single Nvidia RTX 3090 GPU.
@@ A.1 Experimental setting
Table 4: Hyperparameters regarding the Graph-JEPA design for the TUD datasets.

### v2

@@ 3.1 Spatial Partitioning
We base the initial part of our architecture on the recent work of (He et al., 2023), but similar ideas consisting of graph partitioning have been proposed before for Graph SSL (Jin et al., 2020). This step consists of creating different subgraphs (patches), similar to how Vision Transformers (ViT) (Dosovitskiy et al., 2020) operate on images. We rely on the METIS (Karypis & Kumar, 1998) graph clustering algorithm, which partitions a graph into a pre-defined, non-overlapping number of clusters $p$ (Fig. 2a), such that the number of within-cluster links is much higher than between-cluster links. Note that having non-overlapping subgraphs can be problematic since edges can be lost in this procedure, and it is possible to end up with empty "subgraphs". In order to maintain a simple notion of locality in each subgraph and avoid completely empty subgraphs, a one-hop neighborhood expansion [^1] Connecting all (not already present) adjacent nodes from the input graph with each node present in the subgraph. of the nodes in each extracted subgraph is performed (Fig. 2b).
@@ 3.2 Subgraph Embedding
After partitioning the graph, we learn a representation for each subgraph through a GNN (Fig. 2c.). The specific choice of the GNN is arbitrary and depends on what properties one wishes to induce in the representation. The learned node embeddings are mean pooled to create a vector representation for each subgraph: $\{h_{1}...h_{p}\},h\in\mathbb{R}^{d}$. Given that these embeddings will be used as context or target variables, providing additional information regarding the subgraphs is key in order to guide the predictor network. Otherwise, the prediction task might be too difficult. Thus, we propose to use a positional embedding for each subgraph, which is implemented as the maximum Random Walk Structural Embedding (RWSE) of all the nodes in that subgraph. In this way, the position is characterized consistently for each patch. Formally, a RWSE (Dwivedi et al., 2021) for a node $v$ can be defined as:
$P_{v}=(M_{ii},M^{2}_{ii},\ldots,M^{k}_{ii})$ (1)
with $P_{v}\in\mathbb{R}^{k}$, $M^{k}=(D^{-1}A)^{k}$ the random-walk transition matrix of order $k$, and $i$ is the index of node $v$ in the adjacency matrix. Therefore, $M^{k}_{ii}$ encodes the probability of node $v$ landing to itself after a $k$-step random walk. Given a subgraph $l$, let $V_{l}$ denote the set of nodes in $l$. We define the subgraph RWSE as:
@@
where the $\max$ operation is performed elementwise. A straightforward explanation of the above definition is that the positional information of each subgraph is essentially contained in the node with the highest degree [^2] A more suitable term would be the in-degree, but there is no difference in the undirected case., which will act as an anchor reference when predicting the target subgraphs’ latent representations. Intuitively, knowing $P_{l}$ is useful for the prediction task because the features present in the most representative node will have been diffused the most (via the GNN) into the subgraph.
@@ 3.3 Context and Target Encoding
Given the subgraph representations and their respective positional embeddings, we frame the Graph-JEPA prediction task in a similar manner to I-JEPA (Assran et al., 2023). The goal of the network is to predict the latent embeddings of randomly chosen target subgraphs, given one random context subgraph. The prediction is conditioned on positional information regarding each subgraph. At each training step, we choose one random subgraph as the context $x$ and $m$ others as targets $Y=\{y_{1},\ldots,y_{m}\}$ (Fig. 2d). These subgraphs are processed by the context and target encoders (Fig. 2e) which are parametrized by Transformer encoder blocks (without self-attention for the context) where normalization is applied at first (Xiong et al., 2020). The target encoder uses Hadamard self-attention (He et al., 2023), but other choices, such as the standard self-attention mechanism (Vaswani et al., 2017) are perfectly viable. We can summarize this step as:
$z^{x}=E_{c}(x),\;\\ \\ \;Z^{y}=E_{t}(Y),$ (3)
with $z^{x}\in\mathbb{R}^{d}$ and $Z^{y}\in\mathbb{R}^{m\times d}$. At this stage, we can use the predictor network to directly predict $Z^{y}$ from $z^{x}$. This is the typical formulation of JEPAs, also followed by the popular work of Assran et al. (2023). We argue that learning how to organize concepts for abstract objects such as graphs or networks directly in Euclidean space is suboptimal. In the following subsection, we propose a simple trick to bypass this problem using the encoding and prediction mechanisms in Graph-JEPA. A discussion in Section 4.3 will provide additional insights regarding this topic.
@@ 3.4 Latent Target Prediction
Learning hierarchically consistent concepts (Deco et al., 2021) is considered a crucial aspect of human learning, especially during infancy and young age (Rosenberg & Feigenson, 2013). Networks in the real world often conform to some concept of hierarchy (Moutsinas et al., 2021), and this assumption is frequently used when learning graph-level representations (Ying et al., 2018). Thus, we conjecture that Graph-JEPA should operate in a hyperbolic space, where learned embeddings implicitly organize hierarchical concepts(Nickel & Kiela, 2017; Zhao et al., 2023). This gives rise to another issue: commonly used hyperbolic (Poincaré) embeddings are known to have several tradeoffs between dimensionality and performance(Sala et al., 2018; Guo et al., 2022), which severely limits the expressive ability of the model. Given that graphs can describe very abstract concepts, high expressivity in terms of model parameters is preferred. In simple words, we would ideally like to have a high-dimensional latent code that has a concept of hyperbolicity built into it.
To achieve this, we think of the target embedding as a high-dimensional representation of the hyperbolic angle, which allows us to describe each target patch through its position in the 2D unit hyperbola. Formally, given a target patch $l$, its embedding $Z^{y}_{l}$ and positional encoding $P_{l}$, we express the latent target as:
$\psi^{y}_{l}=\begin{pmatrix}\;cosh(\alpha^{y}_{l})\;\\ \\ \;sinh(\alpha^{y}_{l})\;\end{pmatrix},~{}\alpha^{y}_{l}=\frac{1}{N}\sum_{n=1}^{d}{Z^{y}_{l}}^{(n)},$ (4)
where $cosh(\cdot)$ and $sinh(\cdot)$ are the hyperbolic cosine and sine functions. The predictor module is then tasked with directly locating the target in the unit hyperbola, given the context embedding and the target patch’s positional encoding:
$\hat{\psi}^{y}_{l}=W_{2}(\sigma(W_{1}(z^{x}+P_{l})+b_{1}))+b_{2},$ (5)
where $W_{n}$ and $b_{n}$ represent the n-th weight matrix and bias vector (i.e., n-th fully connected layer), $\sigma$ is an elementwise non-linear activation function, and $\hat{\psi}^{y}_{l}\in\mathbb{R}^{2}$. This allows us to frame the learning procedure as a regression problem, and the whole network can be trained end-to-end (Fig. 2f). In practice, we use the smooth L1 loss as the distance function, as it is less sensitive to outliers compared to the typical L2 loss (Girshick, 2015):
@@
Thus, we are effectively measuring how far away the context and target patches are in the unit hyperbola of the plane, but the targets are actually described through a high dimensional latent code (Eq. 4). We explicitly show the differences between this choice and using the Euclidean or Hyperbolic distances as energy functions (in the latent space) in Section 4.3. Our proposed pretraining objective forces the context encoder to understand the differences in the hyperbolic angle between the target patches, which can be thought of as establishing an implicit hierarchy between them.
Preventing representation collapse.
JEPAs are based on a self-distillation procedure. Therefore, they are by definition susceptible to representation collapse (Assran et al., 2023). This is due to the nature of the learning process, where both the context and target representations have to be learned. We formalize this intuition with an example and argue why there is a need to adopt two well-known training tricks that are prevalent in the literature to prevent representation collapse: i) The stop-gradient operation on the target encoder followed by a momentum update (using an Exponential Moving Average (EMA) of the context encoder weights) (Grill et al., 2020; Chen & He, 2021); ii) a simpler parametrization of the predictor compared to the context and target networks (in terms of parameters)(Chen et al., 2020; Baevski et al., 2022);
Let us simplify the problem through the following assumptions: i) The predictor network is linear; ii) There is a one-to-one correspondence between context and target patches. (This holds also in practice due to Eq. 5); iii) The self-predictive task is a least-squares problem in a finite-dimensional vector space. Based on these assumptions, we can rewrite the context features as $X\in\mathbb{R}^{n\times d}$, the target coordinates as $Y\in\mathbb{R}^{n\times 2}$, and the weights of the linear model as $W\in\mathbb{R}^{d\times 2}$. The optimal weights of the predictor are given by solving:
@@
Plugging Eq. 8 into Eq. 7 and factorizing $Y$, the least squares solution leads to the error:
$\left\|(X(X^{T}X)^{-1}X^{T}-I_{n})Y\right\|^{2}$ (9)
Thus, the optimality of a linear predictor is defined by the orthogonal projection of $Y$ onto the orthogonal complement of a subspace of $Col(X)$. As is commonly understood, this translates to finding the linear combination of $X$ that is closest, in terms of $\left\|\cdot\right\|^{2}$, to $Y$. Similarly to what was shown by Richemond et al. (2023), we argue that this behavior unveils a key intuition: The target encoder, which estimates $Y$ must not share weights or be optimized via the same optimizer as the context encoder. If that were the case, the easiest solution to Eq. 9 would be learning a representation that is orthogonal to itself, i.e., the $\mathbf{0}$ vector, leading to representation collapse. Using a well-parametrized EMA update is what allows us to bypass this problem. In practice, even with the slower dynamics induced by the EMA procedure, it is possible to immediately encounter a degenerate solution with a non-linear and highly expressive network. For instance, consider a scenario where the target subgraphs are straightforward and similar. In this case, if the predictor network is sufficiently powerful, it can predict the target representations even without a well-learned context representation. Since the target encoder weights are updated via the EMA procedure, the learned representations will tend to be uninformative. Therefore, implementing the predictor network as a simpler and less expressive network is crucial to achieving the desired training dynamics.
@@ 4 Experiments
The experimental section introduces the empirical evaluation of the Graph-JEPA model in terms of downstream performance on different graph datasets and tasks, along with additional studies on the latent space’s structure and the encoders’ parametrization. Furthermore, a series of ablation studies are presented in order to elucidate the design choices behind Graph-JEPA.
@@ 4.1 Experimental setting
We use the TUD datasets (Morris et al., 2020) as commonly done for graph-level SSL (Suresh et al., 2021; Tan et al., 2023). We utilize seven different graph-classification datasets: PROTEINS, MUTAG, DD, REDDIT-BINARY, REDDIT-MULTI-5K, IMDB-BINARY, and IMDB-MULTI. We report the accuracy of ten-fold cross-validation for all classification experiments over five runs (with different seeds). It is worth noting that we retrain the Graph-JEPA model for each fold without ever having access to the testing partition in both the pretraining and fine-tuning stages. We use the ZINC dataset for graph regression and report the Mean Squared Error (MSE) over ten runs (with different seeds), given that the testing partition is already separated. To produce the unique graph-level representations, we feed all the subgraphs through the trained target encoder and then use mean pooling, obtaining a single feature vector $z_{G}\in\mathbb{R}^{d}$ that represents the whole graph. This high-dimensional is then used to fit a linear model with L2 regularization for the downstream task. Specifically, we employ Logistic Regression with L2 regularization on the classification datasets and Ridge Regression for the ZINC dataset. For the datasets that do not natively have node and edge features, we use a simple constant (0) initialization. The subgraph embedding GNN (Figure 2c.) consists of the GIN operator with support for edge features Hu et al. (2019), often referred to as GINE. The neural network modules were trained using the Adam optimizer Kingma & Ba (2014) and implemented using PyTorch Paszke et al. (2019) and PyTorch-Geometric Fey & Lenssen (2019), while the linear classifiers and cross-validation procedure were implemented through the Scikit-Learn library Pedregosa et al. (2011). All experiments were performed on Nvidia RTX 3090 GPUs. Finally, 1 shows the JEPA-specific hyperparameters used for the following experiments.
@@
Table 1: Values of Graph-JEPA specific hyperparameters for the experiments on the TUD datasets.